Chapter 11, Problem 004

Neutron stars consist only of neutrons and have unbelievably high densities. A typical mass and radius for a neutron star might be 2.0 x 1028 kg and 1.1 x 103 m. (a) Find the density of such a star. (b) If a dime (V = 2.0 x 10-7 m3) were made from this material, how much would it weight (in pounds)?

Chapter 11, Problem 091

The karat is a dimensionless unit that is used to indicate the proportion of gold in a gold-containing alloy. An alloy that is one karat gold contains a weight of pure gold that is one part in twenty-four. What is the volume of gold in a 14.0-karat gold necklace whose weight is 2.00 N?
Answer #1

problem 004)

a)
density = mass/volume

= 2*10^28/(4/3*pi*(1.1*10^3)^3)

= 3.6*10^18 kg/m^3 <<<<<<<--------------------Answer

b) mass of the dime, m = density*volume

= 3.6*10^18*2*10^-7

= 7.20*10^11 kg

weight = 7.2*10^11*2.2 (since 1 kg = 2.2 pounds)

= 1.6*10^13 pounds <<<<<<<--------------------Answer

problem 091)

mass of the alloy, m = W/g

= 2/9.8

= 0.20408 kg

= 204.08 grams

mass of the gold in the alloy, m_gold = (14/24)*204.08

= 119 grams (or) 0.119 kg

volume of gold, V = mass/density

= 119/19.32

= 6.16 cm^3 (or) 6.16*10^-6 m^3 <<<<<<<--------------------Answer